Shad is 3 ft. from a lamppost that is 12 ft. high. If Shad is 5 1/2 ft. tall, How long is Shad's shadow?
Answer by RAY100(1637)

A rough sketch helps to visualize.
.
Large rt triangle, square corner on right, height an right is 12, other leg, horizontal, is x.
.
Smaller triangle is drawn inside , with height 5.5.,,,horizontal leg is (x-3).
.
horizontal difference between heights is 3
.
Problem is proportional triangles, one 12 height by (x) leg,,,other 5.5 high,,(x-3) leg
.
12/x = 5.5 / (x-3)
.
cross multiply
.
12x -36 =5.5 x
.
6.5 x = 36
.
x=5.54,,,,lamppost shadow lg
.
Therefore Shad's shadow lg = (x-3) = 2.54
.
checking
.
12/5.54 = 5.5 /2.54,,,,,ok